OpenStudy (anonymous):

What percent of 90 is 100? A. 90% B. 1.11% C. 111% D. 9%

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

or D

OpenStudy (anonymous):

90% And yeah, it's pretty much A. :p

OpenStudy (anonymous):

It's just mainly 9/10 or 90/100.

OpenStudy (anonymous):

:) okay thank you

OpenStudy (anonymous):

What percent of 35 is 21? A. 0.60 B. 60% C. 31% D. 167%

OpenStudy (anonymous):

B is what i think

OpenStudy (anonymous):

@kaylynn_013 if the question asks you to simply the fraction then it's 9%, but they didn't ask.

OpenStudy (compassionate):

Kylynn, to get the percentage of a number simply multiple that number by the decimal.

OpenStudy (anonymous):

100 is (100/90) percent of 90. 100/90 = 1.111 = 111%

OpenStudy (anonymous):

x/100 %of 35 =21 ie: x/100 * 35 =21 x=21/35 *100 x = 60%

OpenStudy (anonymous):

wow i got all confused

OpenStudy (anonymous):